OpenStudy (anonymous):

solve for t 50= -16T^2 +70T + 3

OpenStudy (anonymous):

round your answer to the nearest hundredth

OpenStudy (anonymous):

first you might want to simplify the quadratic equation by subtracting both sides by 50 then you can apply the quadratic formula to solve for t \(\Large \begin{array}{*{20}c} {T= \frac{{ - \color{blue}{b} \pm \sqrt {\color{blue}{b}^2 - 4 \color{red}{a} \color{green}{c}} }}{{2\color{red}{a}}}} & {{\rm{when}}} & {\color{red}{a}T^2 + \color{blue}{b}T +\color{green}{c} = 0} \\ \end{array}\)

OpenStudy (anonymous):

so just plug it in to the quadratic formula?

OpenStudy (anonymous):

yes after you simplify the equation

OpenStudy (anonymous):

you should have two solutions

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

are you solving for it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

- 70 +- 70 sqrt 192 / 32